OpenStudy (anonymous):

ln(7x)=1 SOLVE FOR X

OpenStudy (anonymous):

\[\ln ( 7x) = 1\]do this step and tell me what u get\[\huge e^{\ln(7x)} = e^1\]

OpenStudy (anonymous):

tbh Im confused on this simple equation...

OpenStudy (anonymous):

hmm...I'm not so sure I know how to do that either

OpenStudy (anonymous):

\[\huge e^{\ln (7x)} = e^1\]\[\huge 7x = e\]

OpenStudy (anonymous):

@Brad1996 and @megachomehead6 Both of you need to catch up with foundation :(

OpenStudy (anonymous):

Basic rule: Natural log and exponent e cancels out :)

OpenStudy (anonymous):

Oh okay let me see if i get the correct answer now

OpenStudy (anonymous):

do you know how to solve\[\log_e7x=1\]

OpenStudy (anonymous):

Nope, the reason I don't know how to do any of this is because my professor assigned multiple sections over the weekend and didn't teach us how to do any of it... thanks for trying though